OpenStudy (sarahc):

Find an equation in standard form for the ellipse with the vertical major axis of length 16, and minor axis of length 4. x squared divided by 64 plus y squared divided by 4 = 1 x squared divided by 2 plus y squared divided by 8 = 1 x squared divided by 8 plus y squared divided by 2 = 1 x squared divided by 4 plus y squared divided by 64 = 1

OpenStudy (sarahc):

@jim_thompson5910

OpenStudy (tkhunny):

It's an ellipse. \(\dfrac{x^{2}}{}+\dfrac{y^{2}}{} = 1\) Vertical Major axis length of 16. 16/2 = 8 So far: \(\dfrac{x^{2}}{}+\dfrac{y^{2}}{8^{2}} = 1\) minor axis of length 4. 4/2 = 2 \(\dfrac{x^{2}}{2^{2}}+\dfrac{y^{2}}{8^{2}} = 1\) Jut one piece at a time. No need to be overwhelmed by the whole thing at once. Okay, that was your free one. Now, you do one.